Tyler is taking piano lessons. Because he doesn't practice as often as his parents would like him to, he has to pay for part of the lessons himself. His parents pay him $0.48 to do the laundry and $1.28 to mow the lawn. In one month, he does the laundry 6 more times than he mows the lawn. If his parents pay him $16.96 that month, how many times did he mow the lawn?
Answer by checkley77(12844)

1.28x+.48(x+6)=16.96
1.28x+.48x+2.88=16.96
1.76x=16.96-2.88
1.76x=14.08
x=14.08/1.76
x=8 is the number of times he mowed the lawn.
Proof:
1.28*8+.48(8+6)=16.96
10.24+.48*14=16.96
10.24+6.72=16.96
16.96=16.96